OpenStudy (anonymous):

I need some serious help. If you reply, can you type the question number you are solving? Thanks Sal's Sandwich Shop sells wraps and sandwiches as part of its lunch specials. The profit on every sandwich is $2 and the profit on every wrap is $3. Sal made a profit of $1,470 from lunch specials last month. The equation 2x + 3y = 1,470 represents Sal's profits last month, where x is the number of sandwich lunch specials sold and y is the number of wrap lunch specials sold.

OpenStudy (anonymous):

2. Describe how you would graph this line using the slope-intercept method. Be sure to write using complete sentences. 3. Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. 4. Graph the function. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work or you may use graphing technology. 5. Suppose Sal's total profit on lunch specials for the next month is $1,593. The profit amounts are the same: $2 for each sandwich and $3 for each wrap. In a paragraph of at least three complete sentences, explain how the graphs of the functions for the two months are similar and how they are different.

OpenStudy (anonymous):

6. Below is a graph that represents the total profits for a third month. Write the equation of the line that represents this graph. Show your work or explain how you determined the equations.

OpenStudy (amoodarya):

you have two points (400,0) ,(0,300) first slope =\[\frac{ y_2-y_1 }{ x_2-x_1 }=\frac{0-300}{400-0}=\frac{-3}{4}\\\]

OpenStudy (anonymous):

For which question is this?

OpenStudy (amoodarya):

then a line with slope m and passing point (x1,y1) is y-y1=m(x-x1) \[y-y_1=m(x-x_1)\\(400,0)\\m=\frac{-3}{4}\\y-0=\frac{-3}{4}(x-400)\]

OpenStudy (aaronandyson):

Find y intercept by letting x = 0 3y = 1470

OpenStudy (anonymous):

Gah! I'm confused! Which question are you answering?

OpenStudy (aaronandyson):

Okay. First find y

OpenStudy (anonymous):

In question 2?

OpenStudy (aaronandyson):

next you need to find the slope of the line 2x+3y = 1470 so 3y = -2x + 1470 then divide by 3 and you get \[y = -\frac{ 2 }{ 3 } + 490\]

OpenStudy (aaronandyson):

@TrinPlays

OpenStudy (anonymous):

I'm just confused. Is this turning the equation into slope intercept form?

OpenStudy (aaronandyson):

Question 2. Yes.

OpenStudy (anonymous):

Oh okay. Thanks for clearing that up.

OpenStudy (aaronandyson):

function notation \[f(x) =- \frac{ 2 }{ 3 } + 490\]

OpenStudy (amoodarya):

i wrote for the last attachment and in form of intercept is \[\frac{x}{400}+\frac{y}{300}=1\]

OpenStudy (aaronandyson):

This means for every 30 sandwiches sold 20 wraps decrease.

OpenStudy (aaronandyson):

Now Graph this function by lettin y = 0 2x = 1470

OpenStudy (aaronandyson):

x = 735

OpenStudy (anonymous):

Ugh so confused

OpenStudy (anonymous):

yoooooooooooooooooooooooo

OpenStudy (anonymous):

FLVS is a pain